# Development of a comprehensive PMT optical model for the JUNO experiment

### v1

@@ Abstract
There are 17,612 20-inch photomultiplier tubes (PMTs) installed at the Jiangmen Underground Neutrino Observatory (JUNO). Developing a precise optical model for the PMTs is crucial for enhancing the accuracy of detector simulations and refining the energy response model at JUNO. In this study, we established a comprehensive PMT optical model based on prior studies, taking into account the non-uniformity of photon detection efficiency (PDE) across the PMT surface and the variances in PDE as well as reflections among different PMTs. By collecting reflectance data from 669 PMTs and utilizing PDE data from mass testing systems, we estimated the thickness maps of the photocathode (PC) and the anti-reflective coating (ARC) for each PMT. We also determined the collection efficiency (CE) by decomposing PDE with consideration of the optical processes occurring within the PMTs. The refractive index and extinction coefficient of both the PC and ARC, along with the escape factor, were evaluated over a broad wavelength range of 300 nm to 700 nm, covering the entire spectrum of interest for JUNO. Compared to the prediction from a simplified PMT optical model, which assumes uniform PC and ARC across all PMTs of the same type, the further developed PMT optical model yields much more reflectance for HPK PMTs and less for NNVT PMTs, and the change in PDE is at the level of a few percent. This comprehensive PMT optical model also provides a valuable reference for other PMT-based applications.
@@ 1 Introduction
The photomultiplier tube (PMT) has served as an optimal photosensor for numerous neutrino projects over the past few decades and remains pivotal in current and future initiatives. One such project is the Jiangmen Underground Neutrino Observatory (JUNO), which has deployed 17,612 20-inch PMTs and 25,600 3-inch PMTs in its central detector. The spherical arrangement of these PMTs achieves up to 75% photocathode (PC) coverage. Among the 20-inch PMTs, about 5,000 are dynode PMTs manufactured by Hamamatsu Photonics K.K. (HPK), while the remainder are MCP-PMTs produced by Northern Night Vision Technology Co. (NNVT) [9]. These PMTs are used to capture light from the 20 kton liquid scintillator contained within an acrylic sphere with a diameter of 35.4 m. The detailed design of the JUNO detector is described in [5]. JUNO’s primary goal is to determine the neutrino mass ordering and precisely measure the neutrino oscillation parameters by analyzing oscillation patterns within the neutrino energy spectrum from the Taishan and Yangjiang nuclear power plants, located approximately 53 km from the JUNO site. Additionally, JUNO has an extensive physics program focused on detecting neutrinos from various natural sources, as outlined in [4, 7].
@@
A precise understanding of the detector response is crucial for determining the neutrino mass ordering and conducting other physics analyses in JUNO [2]. Specifically, the detector response primarily involves the energy resolution, energy non-uniformity, and energy non-linearity within the JUNO detector. The responses of the PMTs, particularly the photon detection efficiency (PDE), significantly influence these energy responses and must be carefully managed. Understanding the PDE of PMTs is complex, as it depends on various factors, including the working media, the angle of incidence (AOI) of photons, photon energy, positions where the light hits the PC, and the shape as well as the internal structure of PMTs. The influence of these factors on the PDE is discussed in detail in [8]. Within the JUNO experiment, the PDE of 20-inch PMTs has been thoroughly characterized via the PMT mass testing system, including the average PDE measured by the container system for all PMTs and the PDE uniformity measured by the scanning station system for a subset, with results summarized in [1]. However, the measured PDE cannot be directly used in detector simulations to predict energy responses because the PMTs operate in water within the JUNO detector, rather than in air, and the AOIs, energies, and positions of photons hitting the PMTs differ from those in the mass testing system. In fact, these factors can vary significantly from event to event in the JUNO detector and lead to different PDE. Therefore, a PMT optical model is essential to accommodate all these factors while adhering to constraints from the mass testing data.
@@
A new optical model for PMTs has been proposed in [8], which accounts for the PMT window as a composite structure comprising the glass substrate, a thin anti-reflective coating (ARC), and a PC. The key parameters of this model were determined from detailed measurements performed on three 20-inch PMTs: two from NNVT and one from HPK. By incorporating the optical properties of the PC and ARC obtained from these measurements into the JUNO simulation program, the energy response of the JUNO detector has been evaluated and predicted under the assumption that PMTs of the same type with the identical optical properties of PC and ARC [3]. However, variations in PDE exist among PMTs, and the properties of the PC and ARC could be non-uniform. Moreover, PMT reflectance exhibits significant variation, which must also be taken into account, as light reflected from the PMTs has a substantial impact on the light yield and energy response in the JUNO detector. In this work, these effects are incorporated by extending the PMT optical model originally developed in [8]. The extension utilizes PDE data obtained from both the container system and the scanning station system, as well as reflection data collected in this study. The resulting comprehensive PMT optical model enables the decoupling of quantum efficiency (QE) and collection efficiency (CE) from the measured PDE, while accounting for the non-uniformity of PC and ARC. Specifically, QE is determined based on the photon absorption ratio in the PC and the escape factor, which represents the fraction of photoelectrons that successfully escape from the PC. Using this model, we derived thickness maps of the PC and ARC for all 17,612 20-inch PMTs, along with their corresponding CE values. The model covers the entire range of photon wavelengths relevant to JUNO, from 300 nm to 700 nm. This comprehensive approach also provides valuable insights for other PMT-based projects seeking to improve their detector performance.
@@
The subsequent sections are organized as follows: we first review the PMT optical model developed in [8], followed by a detailed discussion of the comprehensive PMT optical model. We then introduce the PDE data used in this study, and illustrate the method employed for the PMT reflection measurements and report the reflectance results. Subsequently, the extracted basic parameters for the PMT optical simulation are discussed, including the thickness maps of PC and ARC and CE for individual PMTs, as well as the refractive index and extinction coefficient over an extended wavelength range for three types of PMTs. In the end, the comparison of the average reflectance and PDE calculated from the simplified and comprehensive models is presented.
@@ 2.1 Recap of the PMT optical model
The PDE could be denoted as a function of the photon wavelength ($\lambda$), the AOI ($\alpha$), and the position ($\vec{r}$) of the photon hitting the PC of the PMTs. Specifically, the measured PDE can be contributed by multiple positions in addition to the point of incidence on the PC, as a fraction of the incident light could penetrate the PC and propagate within the PMTs. Consequently, the PDE can be expressed as shown in Eq. 2.1 and is discussed in detail in [8]:
$\mbox{PDE}(\lambda,\alpha,\vec{r})=\sum_{j}a_{j}(\lambda,\alpha,\vec{r})\cdot\beta_{j}\cdot\rho_{j}(\lambda,\vec{r})\cdot\mbox{CE}_{j}(\vec{r})=\sum_{j}a_{j}(\lambda,\alpha,\vec{r})\cdot\beta_{j}\cdot F_{j}(\lambda,\vec{r})$ (2.1)
Here, $j$ represents the $j$-th detected photon, and $a_{j}$ represents the absorption ratio of the incident light, and $\rho_{j}$ is the escape factor. The product of $\rho$ and CE is denoted as an $F$-factor. The coefficient $\beta$ weights the absorption from different positions that contribute to the measured PDE. In an ideally uniform PC, the positional dependencies of $a$ and $\rho$ would be negligible. While this assumption may hold in certain applications, it is generally not valid for large PMTs, such as those employed in JUNO.
The CE is determined by the electric field within the PMTs, which typically varies with the positions on the PC. By selecting several reference positions with known $F$-factors on the PC, the $F$-factor at any other position can be derived through interpolation. Therefore, Eq. 2.1 can be rewritten as:
$\mbox{PDE}=\sum_{j}a_{j}\sum_{k=1}^{n}\beta_{jk}F_{k}$ (2.2)
where $\beta$ represents the weight coefficient, and $F_{k}$ denotes the $F$-factor at the $k$-th reference point. A procedure for deriving the $F$-factors at these reference points has been proposed in [8]. The coefficient $\beta$ can be obtained by simulating the optical processes inside the PMTs using a simulation toolkit. The absorptance $a$, reflectance, and transmittance of incident photons on the PC can be calculated based on the optical theory, using the known optical properties of the PC and ARC, including the refractive index ($n$), extinction coefficient ($k$), and thickness ($d$). In previous work, reflectivity data collected by immersing PMTs in the linear alkylbenzene are used to extract values for $n$, $k$, and $d$ for two NNVT PMTs and one HPK PMT across a wavelength range of 390 nm to 500 nm. These PMTs represent three different types of 20-inch PMTs utilized in the JUNO detector. For NNVT PMTs, there are two types known as NNVT normal-QE PMT and NNVT high-QE PMT, which was produced during the early stages of PMT mass production and manufactured with improved QE achieved through optimized PC technology, respectively.
@@ 2.2 Extension of the PMT optical model
The following assumptions have been made to enable the PMT optical model to account for PDE non-uniformity within a single PMT and PDE variations among different PMTs:
@@
  • PMTs of the same type share identical spectral responses for the refractive index and extinction coefficient of both the PC and the ARC. The values of $n$ and $k$ obtained in [8], as well as the escape factor $\rho$, are employed in this study for the three types of PMTs, covering wavelengths from 390 nm to 500 nm. Utilizing dispersion relations and available data in the literature, we extend this range down to 300 nm and up to 700 nm, encompassing the relevant spectrum for JUNO.
@@
  • Variations in QE and reflectance for a single PMT are induced by non-uniform thicknesses of the PC and ARC. The thickness map is determined in this work using Eq. 2.2, with inputs derived from PDE and reflectance data for each 20-inch PMT. These inputs will be discussed in detail in Sec. 3.
@@
  • Due to the absence of specific QE information, decoupling the position dependence of QE and CE from the measured PDE exists some problems. Therefore, we assume that the CE remains constant in the PC region for zenith angles smaller than 50 degrees (in regions far from the equator) when calculating the thickness maps of the PC and ARC. Thereafter, the positional dependence of CE and its variations among different PMTs are incorporated into Eq. 2.2.
@@
  • For simplification, we assume that both QE and CE are independent of azimuthal angles on the PC. This assumption is a reasonable approximation for the QE of both NNVT PMTs and HPK PMTs, as demonstrated in [1]. Regarding CE, while it is azimuthal angle-dependent for HPK PMTs near the equator due to the orientation of the focusing electrode, the random arrangement of these electrodes during PMT installation in the JUNO detector largely mitigates their impact on energy responses, as contributions from multiple PMTs during physical events sufficiently average out these variations.
@@
Given the assumptions above, the thickness map of the PC and ARC, along with the CE at small zenith angles, can be derived by simultaneously fitting the PDE and reflectance data using Eq. 4.1 for all 20-inch PMTs. To model the position dependence of the thickness for both the PC and the ARC, an empirical formula is employed:
$d_{PC,ARC}=c_{2}\times\theta^{2}+c_{1}\times\theta+c_{0}$ (2.3)
@@
where $\theta$ represents the zenith angle of the position on the PC, while $c_{0}$, $c_{1}$, and $c_{2}$ are free parameters determined through the fitting process. Fig. 1 illustrates the contours of reflectance (top) and absorptance (bottom) at 415 nm and 420 nm, respectively, calculated as a function of the thicknesses of the ARC (x-axis) and PC (y-axis) for the HPK PMT (left), NNVT normal-QE PMT (middle), and NNVT high-QE PMT (right). The refractive index and extinction coefficient of the ARC and PC reported in [8] are used in these calculations. Notably, the different types of PMTs exhibit distinct dependencies in their thickness maps. The fitting procedure aims to identify the optimal path on the contours that aligns closely with both the reflection and absorption data, the latter of which is inherently incorporated into the PDE data. Once the thickness maps of the PC and ARC are established, absorption ratios can be calculated for any position on the PC. The weight coefficients are obtained by simulating the optical processes within the PMTs using a simulation program developed based on Geant4. Subsequently, the CE at the reference points can be determined using the known PDE values at those points via Eq. 2.2. As a result, a complete CE curve can be constructed for each PMT.
@@
Figure 1: Contours of reflectance (top) and absorptance (bottom) at 415 nm and 420 nm, respectively, as a function of thicknesses of the ARC (x-axis) and PC (y-axis) for the HPK PMT (left), NNVT normal-QE PMT (middle) and NNVT high-QE PMT (right).
@@ 3 PDE and reflectance data
In this work, both the PDE results measured by the mass testing systems including the container system and the scanning station system and the newly collected reflectance data are utilized to constrain the thicknesses of the PC and ARC.

### v2

@@ Abstract
There are 17,612 20-inch photomultiplier tubes (PMTs) installed at the Jiangmen Underground Neutrino Observatory (JUNO). Developing a precise optical model for the PMTs is crucial for enhancing the accuracy of detector simulations and refining the energy response model at JUNO. In this study, we established a comprehensive PMT optical model based on the prior studies, taking into account the non-uniformity of the photon detection efficiency (PDE) across the PMT surface and the variances in the PDE as well as reflectance among the different PMTs. By collecting the reflectance data from 669 PMTs and utilizing the PDE data from the mass testing systems, we estimated the thickness maps of the photocathode (PC) and the anti-reflective coating (ARC) for each PMT. We also determined the collection efficiency (CE) by decomposing the PDE with consideration of the optical processes occurring within the PMTs. The refractive index and extinction coefficient of both the PC and ARC, along with the escape factor, were evaluated over a broad wavelength range of 300 nm to 700 nm, covering the entire spectrum of interest for JUNO. Compared to the prediction from a simplified PMT optical model, which assumes the uniform PC and ARC across all the PMTs of the same type, the further developed PMT optical model yields much more reflectance for the HPK PMTs and less for the NNVT PMTs, and the change in the PDE is at the level of a few percent. This comprehensive PMT optical model also provides a valuable reference for other PMT-based applications.
@@ 1 Introduction
The photomultiplier tube (PMT) has served as an optimal photosensor for numerous neutrino projects over the past few decades and remains pivotal in current and future initiatives. One such project is the Jiangmen Underground Neutrino Observatory (JUNO), which has deployed 17,612 20-inch PMTs and 25,600 3-inch PMTs in its central detector. The spherical arrangement of these PMTs achieves up to 75% photocathode (PC) coverage. Among the 20-inch PMTs, about 5,000 are dynode PMTs manufactured by Hamamatsu Photonics K.K. (HPK), while the remainder are MCP-PMTs produced by Northern Night Vision Technology Co. (NNVT) [9]. These PMTs are responsible for capturing light from the 20 kton liquid scintillator contained within an acrylic sphere with a diameter of 35.4 m. The detailed design of the JUNO detector is described in [5]. JUNO’s primary goal is to determine the neutrino mass ordering and precisely measure the neutrino oscillation parameters by analyzing oscillation patterns within the neutrino energy spectrum from the Taishan and Yangjiang nuclear power plants, located approximately 53 km from the JUNO site. Additionally, JUNO has an extensive physics program focused on detecting neutrinos from various natural sources, as outlined in [4, 7].
@@
A precise understanding of the detector response is crucial for determining the neutrino mass ordering and conducting other physics analyses in JUNO [2]. Specifically, the detector response primarily involves the energy resolution, energy non-uniformity, and energy non-linearity within the JUNO detector. The responses of PMTs, particularly the photon detection efficiency (PDE), significantly influence these energy responses and must be carefully managed. Understanding the PDE of PMTs is complex, as it depends on the various factors, including the working media, the angle of incidence (AOI) of photons, photon energy, positions where the light hits the PC, and the shape as well as the internal structure of PMTs. The influence of these factors on the PDE is discussed in detail in [8]. Within the JUNO experiment, the PDE of 20-inch PMTs has been thoroughly characterized via the PMT mass testing system, including the average PDE measured by the container system for all the PMTs and the PDE uniformity measured by the scanning station system for a subset, with results summarized in [1]. However, the measured PDE cannot be directly utilized in detector simulations to predict energy responses because the PMTs operate in water within the JUNO detector, rather than in air, and the AOIs, energies, and positions of photons hitting the PMTs differ from those in the mass testing system. In fact, these factors can vary significantly from event to event in the JUNO detector and lead to the different PDE. Therefore, a PMT optical model is essential to accommodate all these factors while adhering to constraints from the mass testing data.
@@
A new optical model for the PMTs has been proposed in [8], which accounts for the PMT window as a composite structure comprising the glass substrate, a thin anti-reflective coating (ARC), and a PC. The key parameters of this model were determined from detailed measurements performed on three 20-inch PMTs: two from NNVT and one from HPK. By incorporating the optical properties of the PC and ARC extracted from these measurements into the JUNO simulation program, the energy response of the JUNO detector has been evaluated and predicted under the assumption that PMTs of the same type with the identical optical properties of the PC and ARC [3]. However, the variations in the PDE exist among PMTs, and the properties of the PC and ARC could be non-uniform. Moreover, the PMT reflectance exhibits the significant variations, which must also be taken into account, as light reflected from the PMTs has a substantial impact on the light yield and energy response in the JUNO detector. In this study, these effects are incorporated by extending the PMT optical model originally developed in [8]. The extension utilizes the PDE data acquired from both the container system and the scanning station system, as well as the reflectance data collected in this study. The resulting comprehensive PMT optical model enables the decoupling of the quantum efficiency (QE) and collection efficiency (CE) from the measured PDE, while accounting for the non-uniformity of the PC and ARC. Specifically, the QE is determined based on the photon absorption ratio in the PC and the escape factor, which represents the fraction of photoelectrons that successfully escape from the PC. Utilizing this model, we derived the thickness maps of the PC and ARC for all 17,612 20-inch PMTs, along with their corresponding CE values. The model covers the entire range of the photon wavelengths relevant to JUNO, from 300 nm to 700 nm. This comprehensive approach also provides valuable insights for other PMT-based projects seeking to improve their detector performance.
@@
The subsequent sections are organized as follows: we first review the PMT optical model developed in [8], followed by a detailed discussion of the comprehensive PMT optical model. We then introduce the PDE data utilized in this study, and illustrate the method employed for the PMT reflectance measurements and report the reflectance results. Subsequently, the extracted basic parameters for the PMT optical simulation are discussed, including the thickness maps of the PC and ARC and CE for individual PMTs, as well as the refractive index and extinction coefficient over an extended wavelength range for three types of PMTs. In the end, the comparison of the average reflectance and PDE calculated from the simplified and comprehensive models is presented.
@@ 2.1 Recap of the PMT optical model
The PDE could be denoted as a function of the photon wavelength ($\lambda$), the AOI ($\alpha$), and the position ($\vec{r}$) of the photon hitting the PC of PMTs. Specifically, the measured PDE can be contributed by the multiple positions in addition to the point of incidence on the PC, as a fraction of the incident light could penetrate the PC and propagate within the PMTs. These transmitted photons could hit the PC again and then be absorbed, or even undergo multiple reflections before absorption. Consequently, the PDE can be expressed as shown in Eq. 2.1 and is discussed in detail in [8]:
$\mathrm{PDE}(\lambda,\alpha,\vec{r})=\sum_{j}a_{j}(\lambda,\alpha,\vec{r})\cdot\beta_{j}\cdot\rho_{j}(\lambda,\vec{r})\cdot\mathrm{CE}_{j}(\vec{r})=\sum_{j}a_{j}(\lambda,\alpha,\vec{r})\cdot\beta_{j}\cdot F_{j}(\lambda,\vec{r})$ (2.1)
Here, $j$ represents the $j$-th photon, and $a$ represents the absorption ratio of the incident light, and $\rho$ is the escape factor. The product of $\rho$ and CE is denoted as an $F$-factor. The coefficient $\beta$ weights the absorption from the different positions that contribute to the measured PDE. For the ideally uniform PC and ARC, the positional dependencies of $a$ and $\rho$ would be negligible. While this assumption may hold in certain applications, it is generally not valid for large PMTs, such as those employed in JUNO. The realistic non-uniform optical parameters of the PC and ARC could lead to the non-uniformity of the absorptance, which contributes to the non-uniform PDE.
The CE is determined by the electric field within the PMTs and represents the probability of detecting a photon, which typically varies with the positions on the PC, contributing to the PDE non-uniformity as well. By selecting several reference positions with known $F$-factors on the PC, the $F$-factor at any other position can be derived through interpolation. Therefore, Eq. 2.1 can be rewritten as:
$\mathrm{PDE}=\sum_{j}a_{j}\sum_{k=1}^{n}\beta_{jk}F_{k}$ (2.2)
where $\beta$ represents the weight coefficient, and $F_{k}$ denotes the $F$-factor at the $k$-th reference point. A procedure for deriving the $F$-factors at these reference points has been proposed in [8]. The coefficient $\beta$ can be computed by simulating the optical processes inside the PMTs utilizing a simulation toolkit. The absorptance, reflectance, and transmittance of the incident photons on the PC can be calculated based on the optical theory with the optical properties of the PC and ARC, including the refractive index ($n$), extinction coefficient ($k$), and thickness ($d$). In the previous work, the reflectance data collected by immersing PMTs in the linear alkylbenzene are utilized to extract values for $n$, $k$, and $d$ for two NNVT PMTs and one HPK PMT across a wavelength range of 390 nm to 500 nm. These PMTs represent three different types of 20-inch PMTs equipped in the JUNO detector. For the NNVT PMTs, there are two types known as NNVT normal-QE PMT and NNVT high-QE PMT, which were produced during the early stages of the PMT mass production and manufactured with the improved QE achieved through the optimized PC technology, respectively.
@@ 2.2 Extension of the PMT optical model
The following assumptions have been made to enable the PMT optical model to account for the PDE non-uniformity within a single PMT and the PDE variations among different PMTs:
@@
  • The PMTs of the same type share the identical spectral responses for the refractive index and extinction coefficient of both the PC and ARC. The values of $n$ and $k$, as well as the escape factor, acquired in [8], are employed in this study for the three types of PMTs, covering the wavelengths from 390 nm to 500 nm. Utilizing dispersion relations and available data in the literature, we extend this range down to 300 nm and up to 700 nm, encompassing the relevant spectrum for JUNO.
@@
  • The variations in the QE and reflectance for a single PMT are induced by the non-uniform thicknesses of the PC and ARC. The thickness map is determined in this work according to the optical theory, with inputs derived from the PDE and reflectance data for each 20-inch PMT. These inputs will be discussed in detail in Sec. 3.
@@
  • Due to the absence of the specific QE information, decoupling the position dependence of the QE and CE from the measured PDE exists some problems. Therefore, we assume that the CE remains constant in the PC region for the zenith angles smaller than 70 degrees (in regions far from the equator) when calculating the thickness maps of the PC and ARC. Thereafter, the positional dependence of the CE and its variations among different PMTs are incorporated into Eq. 2.2.
@@
  • For simplification, we assume that both the QE and CE are independent of the azimuthal angles on the PC. This assumption is a reasonable approximation for the QE of both the NNVT PMTs and HPK PMTs, as demonstrated in [1]. Regarding the CE, while it is azimuthal angle-dependent for the HPK PMTs near the equator due to the orientation of the focusing electrode, the random arrangement of these electrodes during the PMT installation in the JUNO detector largely mitigates their impact on energy responses, as the contributions from multiple PMTs during physical events sufficiently average out these variations.
@@
Given the assumptions above, the thickness map of the PC and ARC can be derived by simultaneously fitting the PDE and reflectance data for all the 20-inch PMTs. To model the position dependence of the thickness for both the PC and ARC, an empirical formula is employed:
$d_{\mathrm{PC},\mathrm{ARC}}=c_{2_{\mathrm{PC},\mathrm{ARC}}}\times\theta^{2}+c_{1_{\mathrm{PC},\mathrm{ARC}}}\times\theta+c_{0_{\mathrm{PC},\mathrm{ARC}}}$ (2.3)
@@
where $\theta$ represents the zenith angle of the position on the PC, while $c_{0}$, $c_{1}$, and $c_{2}$ are free parameters determined through the fitting process. Fig. 1 illustrates the contours of the reflectance (top) and absorptance (bottom) at 415 nm and 420 nm, respectively, calculated as a function of the thicknesses of the ARC (x-axis) and PC (y-axis) for the HPK PMT (left), NNVT normal-QE PMT (middle), and NNVT high-QE PMT (right). The refractive index and extinction coefficient of the ARC and PC reported in [8] are used in these calculations. Notably, the different types of PMTs exhibit distinct dependencies in their thickness maps. Once the thickness maps of the PC and ARC are established, the non-uniformity of the absorptance on the PC can be calculated. The weight coefficients are computed by simulating the optical processes within the PMTs utilizing a simulation program developed based on Geant4. Subsequently, the CE at the reference points can be determined with the PDE data at those points via Eq. 2.2, and then the CE at other points can be derived by interpolation.
@@
Figure 1: Contours of the reflectance (top) and absorptance (bottom) at 415 nm and 420 nm, respectively, as a function of the thicknesses of the ARC (x-axis) and PC (y-axis) for the HPK PMT (left), NNVT normal-QE PMT (middle) and NNVT high-QE PMT (right).
@@ 3 PDE and reflectance data
In this study, both the PDE results measured by the mass testing systems including the container system and the scanning station system and the newly collected reflectance data are utilized to constrain the thicknesses of the PC and ARC.